# Liouville polarization and the rigidity of their Lagrangian skeleta in dimension $4$

### v1

@@ 2.3. Liouville polarizations (open case)
A Liouville polarization $({\bf\Sigma},\lambda)$ of $(\Omega,\omega)$ consists of ${\bf\Sigma}=\{(\Sigma_{i},\mu_{i})\}$, where $\Sigma=\cup_{i}\Sigma_{i}\subset\Omega$ is a symplectic divisor as above and $\mu_{i}>0$, together with a Liouville form $\lambda$ on $\Omega\setminus\Sigma$ that is tame along $\Sigma$, with residues $-\mu_{i}$ at $\Sigma_{i}$, and such that the Liouville flow $\phi_{X_{\lambda}}^{t}$ on $\Omega\setminus\Sigma$ is backward complete and “forward complete up to hitting $\Sigma$”: For every point $p\in\Omega\setminus\Sigma$, the flow line starting at $p$ is defined for $t\in(-\infty,t^{+}(p))$, where either $t^{+}(p)=+\infty$ or $\lim_{t\nearrow t^{+}(p)}\phi_{X_{\lambda}}^{t}(p)\in\Sigma$.
The points $p$ with $t^{+}(p)<\infty$ form the basin of attraction ${\mathcal{B}}({\bf\Sigma},\lambda,\Omega)$. The complement of the basin of attraction of $\Sigma$ under the Liouville flow is the skeleton of $(\Omega,\omega,{\bf\Sigma},\lambda)$. This is the set of points whose Liouville flow line is defined for all times. If ${\bf\Sigma}$ has only one component $\Sigma$, we call the polarization smooth, and in analogy with the closed case say that it has degree $d=1/\mu$.
@@
(i) In the context of closed manifolds, we have called polarization the sole weighted divisor ${\bf\Sigma}$, which could be easily separated from the Liouville form. In the context of open manifolds, the situation is more intricate because the homological condition $[{\bf\Sigma}]=\operatorname{PD}([\omega])$ does not make sense anymore. This condition is replaced by the residue and integrability conditions on $\lambda$ and its Liouville flow, making impossible to separate ${\bf\Sigma}$ from $\lambda$ in the above definition.
(ii) Whether or not an exact symplectic manifold $(\Omega,\omega)$ has a Liouville polarization does not seem to be easy to decide, in general. We think that if $\Omega$ is the interior of a compact manifold with smooth boudary, this only holds when $(\Omega,\omega)$ is the interior of a very special kind of Liouville domain $(\overline{\Omega},d\beta)$. Note that for Liouville domains, the Liouville vector field $X_{\beta}$ is transverse to the boundary, pointing outwards. This is in sharp contrast to our Liouville vector fields $X_{\lambda}$, that in view of the integrability condition on $X_{\lambda}$ are “tangent at infinity”.
(iii) The reader may wonder why we insist on assuming $\Omega$ open, and propose a definition that may look a bit artificial. Indeed, if $\Omega$ were assumed compact with smooth boundary, we could as well define our Liouville polarization in terms of a Liouville form on $\Omega\backslash\Sigma$ whose Liouville vector field is tangent to $\partial\Omega$. This makes a perfectly sound definition in this case, that would however discard important examples that we want to consider: open manifolds whose boundaries have corners, like the bidisc and, more generally, symplectic disc bundles over open surfaces as in Example 2.14 below, and $\mathbb{R}^{4}$, which is not compact.
@@
A Liouville polarization $({\bf\Sigma},\lambda)$ of $(\Omega,\omega)$ is extendable if there exists an exact symplectic manifold $(\widehat{\Omega},\widehat{\omega})$ with $\Omega\subset\widehat{\Omega}$ and $\widehat{\omega}|_{\Omega}=\omega$, a symplectic divisor $\widehat{\Sigma}\subset\widehat{\Omega}$ with $\Sigma\subset\widehat{\Sigma}$ and such that the closure $\overline{\Sigma}$ of $\Sigma$ in $\widehat{\Sigma}$ is a compact surface with boundary and $\widehat{\Sigma}$ is a collar neighbourhood of $\overline{\Sigma}$ and a tame Liouville form $\widehat{\lambda}$ on $\widehat{\Omega}\backslash\widehat{\Sigma}$ with $\widehat{\lambda}|_{\widehat{\Omega}\setminus\widehat{\Sigma}}=\lambda|_{\Omega\setminus\Sigma}$ such that the boundary $\partial\Omega$ of $\Omega$ in $\widehat{\Omega}$ is a smooth manifold to which $\widehat{\Sigma}$ is transverse (cf. Figure 2.1). For an extendable polarization $({\bf\Sigma},\lambda)$, we denote $\overline{\Sigma}$ the closure of $\Sigma$ in $\widehat{\Sigma}$. $\diamond$
@@
Figure 2.1. An extension $(\widehat{\Omega},\widehat{\Sigma})$ of $(\Omega,\Sigma)$.
In [27] we give a large class of examples of extendable Liouville polarizations. For the moment, we look at the following model example.
@@
$\textnormal{\small SDB}(\overset{\raisebox{-0.86108pt}[0.0pt][-1.29167pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-3.0mu}{X},\tau,\mu):=\pi^{-1}(\overset{\raisebox{-0.86108pt}[0.0pt][-1.29167pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-3.0mu}{X})\;\subset\;\textnormal{\small SDB}(\Sigma,\tau,c_{1}).$
The basin of attraction is the whole manifold, and the skeleton is empty. Taking an open collar neighbourhood of $X\subset\Sigma$ we see that this Liouville polarization is extendable.
@@
The above construction does not rule out the possibility that different extensions $(\Sigma,\tau,c_{1})$ of $(X,\tau)$ lead to different symplectic manifolds $\textnormal{\small SDB}(\overset{\raisebox{-0.86108pt}[0.0pt][-1.29167pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-3.0mu}{X},\tau,\mu)$. This is not so. While this fact is not crucial for us, its proof is a good warm-up for the main result of this section (Theorem 2.20), and it leads to two lemmas needed later on.
Lemma 2.15.
Let $(X,\tau)$ be a compact symplectic surface with boundary, and let $(\widehat{X},\widehat{\tau})$ be an open collar neighbourhood of $X$. Let $W$ be a neighbourhood of $\widehat{X}$ in $\operatorname{SDB}(\widehat{X},\hat{\tau},\mu)$, and let $\vartheta$ be a bounded closed 1-form on $W\setminus\widehat{X}$. Let $\lambda$ be a Liouville 1-form for $\operatorname{SDB}(\widehat{X},\hat{\tau},\mu)$ as constructed in Example 2.14.
@@
Figure 2.2. The symplectomorphism $f\colon V\to V^{\prime}$ with $f^{*}(\lambda+\vartheta)=\lambda$.
Proof.  This is proven in [24, Lemma A.1] in the case that $\widehat{X}$ is a closed surface. The same proof yields Lemma 2.15. $\square$
@@
The symplectic manifold $\textnormal{\small SDB}(\overset{\raisebox{-0.86108pt}[0.0pt][-1.29167pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-3.0mu}{X},\tau,\mu)$ does not depend on the choice of the extension $(\Sigma,\tau,c_{1})$.
Proof.  For $i=0,1$, let $\bigl{(}\textnormal{\small SDB}_{i}(X,\tau,\mu),\lambda_{i}\bigr{)}$ be defined as above by the inclusion $\iota_{i}\colon(X,\tau)\to(\Sigma_{i},\tau_{i})$. We need to show that $\textnormal{\small SDB}_{0}(\overset{\raisebox{-0.86108pt}[0.0pt][-1.29167pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-3.0mu}{X},\tau,\mu)$ is symplectomorphic to $\textnormal{\small SDB}_{1}(\overset{\raisebox{-0.86108pt}[0.0pt][-1.29167pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-3.0mu}{X},\tau,\mu)$. Let $\widehat{X}_{i}\subset\Sigma_{i}$ be open collar neighbourhoods of $X\subset\Sigma_{i}$. We can assume that there exists a symplectomorphism $\varphi\colon(\widehat{X}_{0},\tau_{0})\to(\widehat{X}_{1},\tau_{1})$ that is the identity on $X$. By the symplectic neighbourhood theorem, there exist open neighbourhoods $U_{i}$ of $\widehat{X}_{i}$ in $\operatorname{SDB}(\widehat{X}_{i},\tau_{i},\mu)$ and a symplectic diffeomorphism $\psi\colon U_{0}\to U_{1}$ that extends $\varphi$.
@@
By Proposition 2.5 (ii) there exists a bounded smooth closed one-form $\vartheta$ on $\operatorname{im}\psi\backslash\widehat{X}_{1}$ such that $\psi_{*}\lambda_{0}-\lambda_{1}=\vartheta$. Let $\widetilde{X}_{1}$ be a closed collar neighbourhood of $X$ that lies in $\widehat{X}_{1}$. Applying Lemma 2.15 (i) with $(X,\widehat{X},W):=(\widetilde{X}_{1},\widehat{X}_{1},\operatorname{im}\psi)$, we obtain open neighbourhoods $V,V^{\prime}$ of $X$ in $\operatorname{im}\psi$ and a $C^{1}$-smooth symplectomorphism $f\colon V\to V^{\prime}$ that equals the identity on $\widehat{X}_{1}$ and takes $\lambda_{1}+\vartheta$ to $\lambda_{1}$. Choose $\varepsilon>0$ so small that $\psi\bigl{(}\textnormal{\small SDB}_{0,\varepsilon}(X,\tau,\mu)\bigr{)}\subset\textnormal{Dom}\,f$. Then the map
@@
where for $x\in X$ we take $t=0$ and for $x\notin X$ we take any $t\geq 0$ such that $\phi_{X_{\lambda_{0}}}^{t}(x)\in\text{Dom}(\varphi)\setminus X$. Then $\Phi(x)$ is well-defined because $\varphi_{*}\lambda_{0}=\lambda_{1}$, and a $C^{1}$-smooth symplectomorphism (see e.g. [24, Section 2.1] if needed). By restriction, we obtain a $C^{1}$-smooth symplectomorphism $\textnormal{\small SDB}_{0}(\overset{\raisebox{-0.86108pt}[0.0pt][-1.29167pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-3.0mu}{X},\tau,\mu)\to\textnormal{\small SDB}_{1}(\overset{\raisebox{-0.86108pt}[0.0pt][-1.29167pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-3.0mu}{X},\tau,\mu)$. The claim in the $C^{\infty}$-category now follows from the Smoothing Lemma 2.21 relegated to the end of this section. $\square$
@@
Let $((\Sigma,\mu),\lambda)$ be a smooth and extendable Liouville polarization of $(\Omega,\omega)$. Then there exists $\varepsilon>0$ and a $C^{1}$-smooth symplectic embedding
@@
$\Phi\colon\textnormal{\small SDB}_{\varepsilon}(\Sigma,\omega|_{\Sigma},\mu)\,\to\,{\mathcal{B}}(\widehat{\Sigma},\hat{\lambda},\widehat{\Omega}),$
where $(\widehat{\Sigma},\hat{\lambda},\widehat{\Omega})$ is an extension of the Liouville polarization $(\Sigma,\lambda)$, with $\Phi|_{\Sigma}=\mbox{id}$ and $\Phi^{*}\widehat{\lambda}=\lambda_{0}$. The map $\Phi$ is constructed by a dynamical conjugacy procedure that guarantees that $\operatorname{im}\Phi\subset{\mathcal{B}}(\Sigma,\widehat{\lambda},\widehat{\Omega})$. Since the Liouville flow associated to $\widehat{\lambda}$ is tangent to $\partial\Omega$ in a neighbourhood of $\widehat{\Sigma}\cap\partial\Omega$, this basin of attraction lies in $\Omega$ and coincides with ${\mathcal{B}}(\Sigma,\lambda,\Omega)$. $\square$
@@
Let $({\bf\Sigma},\lambda)$ be an extendable polarization of an exact symplectic manifold $\Omega$. Let $(\widehat{\Omega},\widehat{\Sigma},\widehat{\lambda})$ be an extension and let $\vartheta$ be a bounded closed $1$-form on a deleted neighbourhood of $\widehat{\Sigma}$ in $\widehat{\Omega}\setminus\widehat{\Sigma}$ that vanishes near $\operatorname{Sing}(\Sigma)$. Then there exists a $C^{1}$-smooth symplectomorphism $\Phi$ between neighbourhoods of $\overline{\Sigma}$ in $\widehat{\Omega}$ such that:
@@
Proof.  For each $i$ choose small open discs $\coprod_{j}D_{i}^{j}$ around the points in $\Sigma_{i}\cap\operatorname{Sing}(\Sigma)$ whose closures are disjoint from the support of $\vartheta$, and take the compact surface $X_{i}:=\widetilde{\Sigma}_{i}\setminus\coprod_{j}D_{i}^{j}$, where $\widetilde{\Sigma}_{i}$ is a closed collar neighbourhood of $\overline{\Sigma}_{i}$ in $\widehat{\Sigma}_{i}$. Then $X:=\coprod_{i}X_{i}$ is a disjoint union, and so is the basin ${\mathcal{B}}(X,\widehat{\lambda})=\sqcup_{i}{\mathcal{B}}(X_{i},\widehat{\lambda})$.
@@
Figure 2.4. ${{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}X}}\subset\widehat{\Sigma}$ and $\operatorname{supp}\vartheta$.
By Lemma 2.17 we can identify a neighbourhood of $X_{i}$ in ${\mathcal{B}}(X_{i},\widehat{\lambda})$ with $\operatorname{SDB}_{\varepsilon_{i}}(X_{i},\widehat{\lambda})$. By Lemma 2.15 (ii) there exists a $C^{1}$-smooth symplectomorphism $\Phi_{i}$ between neighbourhoods of $X_{i}$ in $\widehat{\Omega}$ such that $\Phi_{i}^{*}(\widehat{\lambda}+\vartheta)=\widehat{\lambda}$ and $\Phi_{i}=\mbox{id}$ on $X_{i}$ and near the boundaries of the discs $D_{i}^{j}$. Extend $\Phi_{i}$ by the identity to an open neighbourhood of all of $\widetilde{\Sigma}_{i}$. Then the map $\Phi$ obtained by gluing together the $\Phi_{i}$ is as required. $\square$
Our main interest in the definition of Liouville polarizations is the next result, that allows to use the techniques set up to study the symplectic embedding problems in [24, 25] in the context of exact symplectic manifolds.
@@
Recall that an embedding $\varphi$ is $(\alpha,\alpha^{\prime})$-exact if $\varphi^{*}\alpha^{\prime}=\alpha+df$ for a smooth function $f$.
@@
Let $(\Omega,\omega=d\alpha)$ and $(\Omega^{\prime},\omega^{\prime}=d\alpha^{\prime})$ be exact symplectic manifolds with Liouville polarizations $({\bf\Sigma},\lambda)$ and $({\bf\Sigma}^{\prime},\lambda^{\prime})$, respectively. Assume that $({\bf\Sigma},\lambda)$ is extendable, with extension $(\widehat{\bf\Sigma},\widehat{\lambda})$. Assume further that there exists a symplectic morphism $\phi\colon\widehat{\bf\Sigma}\to{\bf\Sigma}^{\prime}$ that is $(\alpha|_{\Sigma},\alpha^{\prime}|_{\Sigma^{\prime}}$)-exact. Then there exists an $(\alpha,\alpha^{\prime})$-exact symplectic embedding
@@
Proof.  Let $(\widehat{\Omega},\widehat{\Sigma},\widehat{\lambda})$ be the extension of $(\Omega,\Sigma,\lambda)$. Let $\widetilde{\Sigma}$ be a closed collar neighbourhood of $\overline{\Sigma}$ in $\widehat{\Sigma}$. By the symplectic neighbourhood theorem, there exists an open neighbourhood $V$ of $\widetilde{\Sigma}$ in $\widehat{\Omega}$ and a symplectic extension $\psi\colon V\hookrightarrow{\mathcal{B}}({\bf\Sigma}^{\prime},\lambda^{\prime})\subset\Omega^{\prime}$ of $\phi$ that preserves the symplectic plumbing structure (i.e., the Darboux coordinates $(R_{1},\theta_{1},R_{2},\theta_{2})$ near the intersections). Since $\psi^{*}\lambda^{\prime}$ and $\widehat{\lambda}$ are tame and have the same residues,
@@
By Lemma 2.18 there exists a $C^{1}$-smooth symplectomorphism $\Phi_{1}$ between open neighbourhoods of $\overline{\Sigma}$ in ${\mathcal{B}}(\widetilde{\bf\Sigma},\hat{\lambda})$ such that $\Phi_{1}^{*}\widehat{\lambda}=\widehat{\lambda}+\vartheta$ and $\Phi_{1}|_{\widetilde{\Sigma}}=\mbox{id}$. Then the map $\Phi:=\psi\circ\Phi_{1}^{-1}$ provides a symplectic extension of $\phi$ such that $\Phi^{*}\lambda^{\prime}=\widehat{\lambda}$. As in the proof of Proposition 2.16 define
@@
where for $x\in\Sigma$ we take $t=0$ and for $x\notin\Sigma$ we take any $t\geq 0$ such that $\phi_{X_{\lambda}}^{t}(x)\in\text{Dom}(\Phi)\setminus\Sigma$. Then $\Psi_{1}$ is a $C^{1}$-smooth symplectic embedding. It is $(\alpha,\alpha^{\prime})$-exact. Indeed, since ${\mathcal{B}}({\bf\Sigma},\lambda)$ retracts to $\Sigma$, its first homology is generated by cycles in $\Sigma$, on which $\Psi_{1}=\phi$. Since $\phi$ is $(\alpha|_{\Sigma},\alpha^{\prime}|_{\Sigma^{\prime}})$-exact by assumption, $\Psi_{1}$ is $(\alpha,\alpha^{\prime})$-exact. The claim in the smooth category now follows from Lemma 2.21. $\square$
@@
Proof.  It is shown in [32, pp. 831–836] that there exists a $C^{\infty}$-smooth symplectomorphism $\Psi\colon(M,\omega)\to(M^{\prime},\omega^{\prime})$ which is $C^{1}$-close to $\Psi_{1}$. The proof shows that $\Psi$ is $(\alpha,\alpha^{\prime})$-exact whenever $\Psi_{1}$ is $(\alpha,\alpha^{\prime})$-exact: Take Darboux charts $\varphi_{i}\colon B^{4}(a_{i})\to U_{i}\subset M$ such that $\varphi_{i}\bigl{(}B^{4}(a_{i}/2)\bigr{)}=:U_{i}^{\prime}$ form a locally finite covering of $M$. Smoothing a suitable generating function, one replaces $\Psi_{1}$ by $\Psi_{1}^{1}$ that is $C^{\infty}$-smooth on $U_{1}^{\prime}$ and agrees with $\Psi_{1}$ outside $U_{1}$. In fact, convex interpolation of the two generating functions yields a smooth symplectic isotopy $\Psi_{1}^{t}$ from $\Psi_{1}$ to $\Psi_{1}^{1}$ such that $\Psi^{-1}_{1}\circ\Psi_{1}^{t}$ is supported in $U_{1}$. Since $U_{1}=\varphi_{1}(B^{4}(a_{1}))$ is simply-connected, this isotopy is Hamiltonian, and hence $\Psi^{-1}_{1}\circ\Psi_{1}^{1}$ is $(\alpha,\alpha)$-exact, see e.g. [21, Proposition 9.3.1]. Since $\Psi_{1}$ is $(\alpha,\alpha^{\prime})$-exact, we find that $\Psi^{1}_{1}$ is also $(\alpha,\alpha^{\prime})$-exact. In the same way, smoothen $\Psi_{1}^{1}$ to $\Psi_{1}^{2}$ on $U_{2}^{\prime}$, $\dots$, $\Psi_{1}^{k}$ to $\Psi_{1}^{k+1}$ on $U_{k+1}^{\prime}$, $\dots$. The limit map $\Psi$ is then $(\alpha,\alpha^{\prime})$-exact. $\square$
3. Some explicit polarizations and skeleta

### v2

@@ 2.3. Liouville polarisations (open case)
A Liouville polarisation $({\bf\Sigma},\lambda)$ of $(\Omega,\omega)$ consists of ${\bf\Sigma}=\{(\Sigma_{i},\mu_{i})\}$, where $\Sigma=\bigcup_{i}\Sigma_{i}\subset\Omega$ is a symplectic divisor as above and $\mu_{i}>0$, together with a Liouville form $\lambda$ on $\Omega\setminus\Sigma$ that is tame along $\Sigma$, with residues $\mu_{i}$ at $\Sigma_{i}$, and such that the Liouville flow $\phi_{X_{\lambda}}^{t}$ on $\Omega\setminus\Sigma$ is backward complete and “forward complete up to hitting $\Sigma$”: For every point $p\in\Omega\setminus\Sigma$, the flow line starting at $p$ is defined for $t\in(-\infty,t^{+}(p))$, where either $t^{+}(p)=+\infty$ or $\lim_{t\nearrow t^{+}(p)}\phi_{X_{\lambda}}^{t}(p)\in\Sigma$.
The points $p$ with $t^{+}(p)<\infty$ form the basin of attraction ${\mathcal{B}}({\bf\Sigma},\lambda,\Omega)$. The complement of the basin of attraction of $\Sigma$ under the Liouville flow is the skeleton of $(\Omega,\omega,{\bf\Sigma},\lambda)$. This is the set of points whose Liouville flow line is defined for all times. If $\bf\Sigma$} has only one component $\Sigma$, we call the polarisation smooth, and in analogy with the closed case say that it has degree $d=1/\mu$.
@@
(i) In the context of closed manifolds, we have called polarisation the sole weighted divisor $\bf\Sigma$}, which could be easily separated from the Liouville form. In the context of open manifolds, the situation is more intricate because the homological condition ${\bf\Sigma}]=\operatorname{PD}([\omega])$ does not make sense anymore. This condition is replaced by the residue and integrability conditions on $\lambda$ and its Liouville flow, making impossible to separate $\bf\Sigma$} from $\lambda$ in the above definition.
(ii) Whether or not an exact symplectic manifold $(\Omega,\omega)$ has a Liouville polarisation does not seem to be easy to decide, in general. We think that if $\Omega$ is the interior of a compact manifold with smooth boundary, this only holds when $(\Omega,\omega)$ is the interior of a very special kind of Liouville domain $(\overline{\Omega},d\beta)$. Note that for Liouville domains, the Liouville vector field $X_{\beta}$ is transverse to the boundary, pointing outwards. This is in sharp contrast to our Liouville vector fields $X_{\lambda}$, that in view of the integrability condition on $X_{\lambda}$ are “tangent at infinity”.
(iii) The reader may wonder why we insist on assuming $\Omega$ open, and propose a definition that may look a bit artificial. Indeed, if $\Omega$ were assumed compact with smooth boundary, we could as well define our Liouville polarisation in terms of a Liouville form on $\Omega\backslash\Sigma$ whose Liouville vector field is tangent to $\partial\Omega$. This makes a perfectly sound definition in this case, that would however discard important examples that we want to consider: open manifolds whose boundaries have corners, like the bidisc and, more generally, symplectic disc bundles over open surfaces as in Example 2.14 below, and $\mathbb{R}^{4}$, which is not compact.
@@
[…]
@@
[…]
In [29] we give a large class of examples of extendable Liouville polarisations. For the moment, we look at the following model example.
@@
Let $(X,\tau)$ be a compact symplectic surface with smooth non-empty boundary, and let $\mu>0$. Consider an extension $(\Sigma,\tau)$ of $(X,\tau)$ to a closed symplectic surface, whose area ${\mathcal{A}}_{\tau}(\Sigma)$ is an integral multiple $c_{1}$ of $\mu$. Let $\pi\colon\bigl(\textnormal{\small SDB}(\Sigma,\tau,c_{1}),\lambda_{0}\bigr)\to\Sigma$ be the symplectic disc bundle defined in Section 2.1, together with its Liouville form of residue $\mu=-\frac{{\mathcal{A}}_{\tau}(\Sigma)}{c_{1}}$. Then $X_{\lambda_{0}}$ vanishes on the boundary of $\textnormal{\small SDB}(\Sigma,\tau,c_{1})$ by (2.3) and is tangent to the fibers. Hence $\bigl((\overset{\raisebox{-0.60275pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-3.0mu}{X},\mu),\lambda_{0}\bigr)$ is a Liouville polarisation of
$\textnormal{\small SDB}(\overset{\raisebox{-0.60275pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-3.0mu}{X},\tau,\mu):=\pi^{-1}(\overset{\raisebox{-0.60275pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-3.0mu}{X})\;\subset\;\textnormal{\small SDB}(\Sigma,\tau,c_{1}).$
The basin of attraction is the whole manifold, and the skeleton is empty. Taking an open collar neighbourhood of $X\subset\Sigma$ we see that this Liouville polarisation is extendable.
@@
The above construction does not rule out the possibility that different extensions $(\Sigma,\tau,c_{1})$ of $(X,\tau)$ lead to different symplectic manifolds $\textnormal{\small SDB}(\overset{\raisebox{-0.60275pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-3.0mu}{X},\tau,\mu)$. This is not so. While this fact is not crucial for us, its proof is a good warm-up for the main result of this section (Theorem 2.20), and it leads to two lemmas needed later on.
Lemma 2.15.
@@
Figure $2.$ The symplectomorphism $f\colon V\to V^{\prime}$ with $f^{*}(\lambda+\vartheta)=\lambda$.
Proof.  This is proven in [25, Lemma A.1] in the case that $\widehat{X}$ is a closed surface. The same proof yields Lemma 2.15. $\square$
@@
The symplectic manifold $\textnormal{\small SDB}(\overset{\raisebox{-0.60275pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-3.0mu}{X},\tau,\mu)$ does not depend on the choice of the extension $(\Sigma,\tau,c_{1})$.
Proof.  For $i=0,1$, let $\bigl(\textnormal{\small SDB}_{i}(X,\tau,\mu),\lambda_{i}\bigr)$ be defined as above by the inclusion $\iota_{i}\colon(X,\tau)\to(\Sigma_{i},\tau_{i})$. We need to show that $\textnormal{\small SDB}_{0}(\overset{\raisebox{-0.60275pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-3.0mu}{X},\tau,\mu)$ is symplectomorphic to $\textnormal{\small SDB}_{1}(\overset{\raisebox{-0.60275pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-3.0mu}{X},\tau,\mu)$. Let $\widehat{X}_{i}\subset\Sigma_{i}$ be open collar neighbourhoods of $X\subset\Sigma_{i}$. We can assume that there exists a symplectomorphism $\varphi\colon(\widehat{X}_{0},\tau_{0})\to(\widehat{X}_{1},\tau_{1})$ that is the identity on $X$. By the symplectic neighbourhood theorem, there exist open neighbourhoods $U_{i}$ of $\widehat{X}_{i}$ in $\operatorname{SDB}(\widehat{X}_{i},\tau_{i},\mu)$ and a symplectic diffeomorphism $\psi\colon U_{0}\to U_{1}$ that extends $\varphi$.
@@
[…]
@@
where for $x\in X$ we take $t=0$ and for $x\notin X$ we take any $t\geq 0$ such that $\phi_{X_{\lambda_{0}}}^{t}(x)\in\text{Dom}(\varphi)\setminus X$. Then $\Phi(x)$ is well-defined because $\varphi_{*}\lambda_{0}=\lambda_{1}$, and a $C^{1}$-smooth symplectomorphism (see e.g. [25, Section 2.1] if needed). By restriction, we obtain a $C^{1}$-smooth symplectomorphism $\textnormal{\small SDB}_{0}(\overset{\raisebox{-0.60275pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-3.0mu}{X},\tau,\mu)\to\textnormal{\small SDB}_{1}(\overset{\raisebox{-0.60275pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-3.0mu}{X},\tau,\mu)$. The claim in the $C^{\infty}$-category now follows from the Smoothing Lemma 2.21 relegated to the end of this section. $\square$
@@
Let $((\Sigma,\mu),\lambda)$ be a smooth and extendable Liouville polarisation of $(\Omega,\omega)$. Then there exists $\varepsilon>0$ and a $C^{1}$-smooth symplectic embedding
@@
[…]
@@
[…]
@@
[…]
Figure 2.4. $^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}X}}\subset\widehat{\Sigma}$ and $\operatorname{supp}\vartheta$.
[…]
Our main interest in the definition of Liouville polarisations is the next result, that allows to use the techniques set up to study the symplectic embedding problems in [25, 26] in the context of exact symplectic manifolds.
@@
[…]
Then there exists an $(\alpha,\alpha^{\prime})$-exact symplectic embedding
@@
[…]
@@
[…]
@@
where for $x\in\Sigma$ we take $t=0$ and for $x\notin\Sigma$ we take any $t\geq 0$ such that $\phi_{X_{\lambda}}^{t}(x)\in\text{Dom}(\Phi)\setminus\Sigma$. Then $\Psi_{1}$ is a $C^{1}$-smooth symplectic embedding. Since ${\mathcal{B}}({\bf\Sigma},\lambda)$ retracts to $\Sigma$, its first homology is generated by cycles in $\Sigma$, on which $\Psi_{1}=\phi$. Hence, if $\phi$ is $(\alpha|_{\Sigma},\alpha^{\prime}|_{\Sigma^{\prime}})$-exact, $\Psi_{1}$ is $(\alpha,\alpha^{\prime})$-exact. The claim in the smooth category now follows from Lemma 2.21. $\square$
@@
Proof.  It is shown in [35, pp. 831–836] that there exists a $C^{\infty}$-smooth symplectomorphism $\Psi\colon(M,\omega)\to(M^{\prime},\omega^{\prime})$ which is $C^{1}$-close to $\Psi_{1}$. The proof shows that $\Psi$ is $(\alpha,\alpha^{\prime})$-exact whenever $\Psi_{1}$ is $(\alpha,\alpha^{\prime})$-exact: Take Darboux charts $\varphi_{i}\colon B^{4}(a_{i})\to U_{i}\subset M$ such that $\varphi_{i}\bigl(B^{4}(a_{i}/2)\bigr)=:U_{i}^{\prime}$ form a locally finite covering of $M$. Smoothing a suitable generating function, one replaces $\Psi_{1}$ by $\Psi_{1}^{1}$ that is $C^{\infty}$-smooth on $U_{1}^{\prime}$ and agrees with $\Psi_{1}$ outside $U_{1}$. In fact, convex interpolation of the two generating functions yields a smooth symplectic isotopy $\Psi_{1}^{t}$ from $\Psi_{1}$ to $\Psi_{1}^{1}$ such that $\Psi^{-1}_{1}\circ\Psi_{1}^{t}$ is supported in $U_{1}$. Since $U_{1}=\varphi_{1}(B^{4}(a_{1}))$ is simply-connected, this isotopy is Hamiltonian, and hence $\Psi^{-1}_{1}\circ\Psi_{1}^{1}$ is $(\alpha,\alpha)$-exact, see e.g. [22, Proposition 9.3.1]. Since $\Psi_{1}$ is $(\alpha,\alpha^{\prime})$-exact, we find that $\Psi^{1}_{1}$ is also $(\alpha,\alpha^{\prime})$-exact. In the same way, smoothen $\Psi_{1}^{1}$ to $\Psi_{1}^{2}$ on $U_{2}^{\prime}$, $\dots$, $\Psi_{1}^{k}$ to $\Psi_{1}^{k+1}$ on $U_{k+1}^{\prime}$, $\dots$. The limit map $\Psi$ is then $(\alpha,\alpha^{\prime})$-exact. $\square$
3. Some explicit polarisations and skeleta